OpenStudy (anonymous):

which equation represents a line through (-2,1) that is perpendicular to y=-5x+2?

OpenStudy (anonymous):

Well the general equation of a line is y=ax+b, we can find a and b by using these relations: a*-5=-1 -2*a+b=1

OpenStudy (anonymous):

which equation represents a line through (-2,1) that is perpendicular to y=-5x+2?